Tom Adams has received a job offer from a large investment bank as a clerk to an associate banker. His base salary will be $53,000. He will receive his first annual salary payment one year from the day he begins to work. In addition, he will get an immediate $10,000 bonus for joining the company. His salary will grow at 3.3 percent each year. Each year he will receive a bonus equal to 10 percent of his salary. Mr. Adams is expected to work for 25 years.

What is the present value of the offer if the discount rate is 9 percent?
Answer #1

Salary and bonus at year T 1 = 53,000 *1.1 = 58300

Bonus = 10,000

Present value = 10,000 + [ 58300 / ( 0.09 - 0.033) ] * [1 - (1.033/1.09) 25 ] = 765,728.296